A new illness is sweeping the students of SUNY Old Westbury. Termed “Epi-itis”, this disease is characterized by an overwhelming desire to measure the cause, spread, and control of disease. Epi-itis is incurable and associated with an increased risk of death. Epidemiologists conduct a study following 400 students for one semester, beginning September 1st, 2019. On September 1st, 40 students are diagnosed with Epi-itis just from their experience at orientation. From September 2nd through December 31st, an additional 60 students develop Epi-itis. Among all those who developed the illness in the cohort of 400, 10 died at the end of the semester. Based on this information, please answer the following:

QUESTION: Among all those who developed the illness in the cohort of 400, 10 died at the end of the semester. Among the 60 students who developed Epi-it is from September 2nd through December 31st , 20 of them developed the illness on October 1st, 20 on November 1st, and 20 on December 1st. What is the incidence rate of Epi-it is from September 2nd through December 31st? Provide your answer using person-months.

Question 2: Among all those who developed the illness in the cohort of 400, 10 died at the end of the semester. Among the 60 students who developed Epi-it is from September 2nd through December 31st , 20 of them developed the illness on October 1st, 20 on November 1st, and 20 on December 1st.
Answer #1

An incidence rate is the number of new cases of a disease divided by the number of persons at risk for the disease.

Students were removed when they were no longer at risk of the outcome, which was epi-itis. All participants began the study at-risk of developing epi-itis. 20 patients were removed from the denominator of the rate at the beginning of the October month. 20 participants were removed from the denominator of the rate at the beginning of November. 20 participant were removed from the denominator of the rate at the beginning of December. The remaining 300 epi-itis-free patients were removed from the denominator of the rate only at the very end of the semester, which would have been December 31st.

Here, over the semester, on 1st September itself, out of 400, 40 students were diagnosed with Epi-it is. So students not having Epi-itis as on 2nd September is 400-40=360

300 students*4 months= 1200

20students*1months=20

20students*2months=40

20students*3months=60

1200+20+40+60=1320 person months

Incidence rate per person month= 60/1320= 0.0455 per person month (Approx.)